OpenStudy (anonymous):

The quantity p varies inversely as the square of (q + 2). p = 5 when q = 3. Find p when q = 8.

OpenStudy (anonymous):

hint: when y varies inversely as x, it means \[y = \frac{m}{x}\]

OpenStudy (anonymous):

where m is a constant

OpenStudy (anonymous):

so you have to use \[p = \frac{m}{(q+2)^{2}}\] and use the given information to find m

OpenStudy (anonymous):

humm, interesting, that explains well.

OpenStudy (anonymous):

and then use m to find \[p = \frac{m}{(8+2)^2}\]

OpenStudy (anonymous):

Thanks! :)